OpenStudy (anonymous):

Yup, I have NO CLUE where to start. WILL MEDAL AND FAN! A large rectangle has a length of 2 inches and a width of 0.18 inches. A smaller rectangle has a length of x³ inches and a width of 0.49x³ inches. a) What is the area of the large rectangle? b) What is the area of the small rectangle? c) Write a polynomial expression that represents how much greater the area of the larger rectangle is than the area of the smaller rectangle. d) Rewrite this expression as a difference of squares

OpenStudy (campbell_st):

well you must be able to complet (a) for (b) its \[A = x^3 \times 0.49x^3 \] which you should be able to answer

OpenStudy (anonymous):

Yeah I got that part 0.49x^6y=0.36 is the answer to C, how do I figure out D?

OpenStudy (campbell_st):

(c) the difference \[P(x) = 0.49x^6 - 0.36\] (d) \[P(x) = (0.7x^3 - 0.6)(0.7x^3 + 0.6)\]